OpenStudy (accidentalaichan):

Change the given equation into slope-intercept form. -3x+2y=8 Could someone show me the steps of completing this?

OpenStudy (anonymous):

first thing u do is add 3x to both sides so now u have 2y=3x+9 then u divide by 2 and u get y=3/2x+4 and that is ur answer .

OpenStudy (anonymous):

y=mx+b,this the slope intercept form where m=slope, b=y-int 1. transpose x to the right side or add both sides by 3x,simplify 2.then, divide both sides by two, then simplify.. asma already gave you the answer

OpenStudy (accidentalaichan):

Thanks, both of you :)